To test whether all University of Maryland undergraduate students spent more than $300 on textbooks during the spring 2020 semester, on average, D. Pines collected data for a random sample of 76 UMD undergraduate students. If the calculated value for the associated test statistic equals +1.9921 and the population standard deviation is unknown, determine the p-VALUE for the associated test

Group of answer choices

0.05

0.025

0.01

0.0125
Answer #1

We have given,

Sample size =n=76

Degree of freedom =n-1=76-1=75

Right tailed t test .

Population standard deviation not known.

P value =0.025....................by using Excel =TDIST(1.9921,75,1)